OpenStudy (anonymous):

Kelly read 30 pages of her 300-page book in 6 hours. At this rate, how long will it take her to read the entire book?

OpenStudy (anonymous):

\(\bf \Large rate= \frac{no.\ of\ pages}{time} = \frac{30\ pages}{6\ hours}= \color{red}{??} page\ per\ hr\\ \Large \color{red}{??} page\ per\ hr=\frac{300\ pages}{ \color{green}{x}\ time} \) nos solve for \(\color{red}{??} \) , then using that solve for \(\color{green}{x}\), which will be your answer